OpenStudy (zidane):

how to solve the triple integral using cartesian coordinates for upper hemisphere of radius R

OpenStudy (anonymous):

\[\int\limits_{-1}^{1}\int\limits_{-\sqrt{1-y^2}}^{\sqrt{1-y^2}}\int\limits_{0}^{\sqrt{1-y^2-x^2}}dzdydx\]

OpenStudy (zidane):

why 1 ? is it suppose to be R? can you solve it? this is not solving the question.

OpenStudy (anonymous):

i am sorry 1 should be r

OpenStudy (anonymous):

ok?

OpenStudy (zidane):

im stuck after i integrate z. help me

OpenStudy (zidane):

help me!

OpenStudy (zidane):

i already got it.